OpenStudy (anonymous):

sin(x) = (1/2)^1/2 between the interval [0, 2pi], I know the answer is sin^-1((1/2)^1/2) but I need the answer in exact fraction form, and I don't know how to do that? What is the exact answer?

OpenStudy (misty1212):

HI!!!\[\sin(x)=\frac{\sqrt2}{2}\] is another way of looking at this question

OpenStudy (misty1212):

look at the unit circle and find the two places where the second coordinate is \(\frac{\sqrt2}{2}\) that will give you the answer

OpenStudy (anonymous):

It still wouldn't work. The answer should have been pi/4 but when I blog it into the answer, it returns wrong. Thank you for trying to help.

OpenStudy (misty1212):

pi over 4, also 3pi over 4 unless it wants degrees for some reason

OpenStudy (anonymous):

thank you, it worked.